Dated 02/01/13 with two new changes and corrections where vector printout does not match postscript file.

I new change: no subscript (calligraphic A) on integral. correction: more space between Tr and Pexp V_{K}(t) = (t + t^{3} − t^{4})(\sqrt{t} + \sqrt{t}^{−1}) V_{K}(e^{2πi/(k+2)}) = \frac{1}{Z} ∫ (Tr Pexp ∮_{K} A) e^{(ik/4π)CS(A)}DA II correction: last subscript is n, not m C_{ijk}η^{kl}C_{lmn} = C_{mjk}η^{kl}C_{lin} III R_{12}R_{23}R_{12} = R_{23}R_{12}R_{23} IV Equations for Lorenz attractor:

\frac{dx}{dt} = σ(y − x)

\frac{dy}{dt} = x(ρ − z) − y

\frac{dz}{dt} = xy − βz

V correction: show difference between v and ν greek nu (v’s have subscripts)

Note: we need to work on graphic ∂_{t}v_{i} + v_{j}∂_{j}v_{i} = −∂_{i}p + ν∂_{j}∂_{j}v_{i} VI new change: bold dx instead of d “ell.” Note: we need to work on the graphic. ∫_{C_{1}} A · dx − ∫_{C_{2}} A · dx = \frac{1}{2π} Φ VII correction: center letter in second term is γ (Greek gamma) and not

r.

correction: last σ is a SUBscript. correction: please find better calligraphic L, more like Palmer.

Supergravity:

L = R − ψ̄_{μ}γ^{μρσ}D_{ρ}ψ_{σ}

r_{S} = 2Gm/c^{2} IX χ = V − E + F 2πχ = ∫_{M} K dA X. Pythagoras:

c^{2} = a^{2} + b^{2}

XI. Babylonian Tablet XII. Golden mean = partial fraction expansion:

lim_{n→∞} \frac{F_{n+1}}{F_{n}} = 1 + \frac{1}{1 + …}

XIII. Archimedes:

v = \frac{2}{3}V

a = \frac{2}{3}A